# On the arithmetic of polynomial ideals

### v1

@@ 4. Arithmetical properties of ${\mathcal{M}{\rm on}}(R)$
Let $I\in{\mathcal{M}{\rm on}}(R)$, and $\mathcal{M}$ a set of monomials in $I$. Then $\mathcal{M}$ generates $I$ if and only if for each monomial $v\in I$ there exists $u\in\mathcal{M}$ such that $u\mid v$.
@@
is an atom of ${\mathcal{M}{\rm on}}(R)$, while back in Section 2 we recalled that it is not an atom of $\mathcal{I}(R)$ (at least when the characteristic of $K$ is not $2$), despite $\{0,1,2,4\}$ being an atom of $\mathcal{P}_{\rm fin}(\mathbb{N})$. Suppose that $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$, with $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$ and $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$. By equation (1) and Remark 4.2, $d+e=4$, $d\geq 1$ and $e\geq 1$. By symmetry, we can also assume $d\leq e$, so the only possibilities for $(d,e)$ are $(1,3)$ and $(2,2)$.
Case $(d,e)=(1,3)$: first, we claim that $Y^{3}\in\mathfrak{b}$. Indeed, if this is not the case, for every monomial $X^{i}Y^{j}\in\mathfrak{b}$ we have either $i\geq 1$ or $j\geq 4$, as $e=3$; the product of such elements with monomials of $\mathfrak{a}$, all of which have positive min-degree, yields elements of the form $X^{k}Y^{l}\in\mathfrak{a}\cdot\mathfrak{b}$, with either $k\geq 1$ or $l\geq 5$. In particular, $Y^{4}\not\in\mathfrak{a}\cdot\mathfrak{b}$, a contradiction. Now suppose $X\notin\mathfrak{a}$; since $d=1$, for every monomial $X^{i}Y^{j}\in\mathfrak{a}$ we have either $i\geq 2$ or $j\geq 1$. The product of such elements with monomials of $\mathfrak{b}$ yields elements of the form $X^{k}Y^{l}$, with either $k+l\geq 5$ or $l\geq 1$. In particular, $X^{4}\not\in\mathfrak{a}\cdot\mathfrak{b}$, again a contradiction; therefore, $X\in\mathfrak{a}$. We then obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, absurd by Lemma 4.1, as none of the monomials $X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}$ divide $XY^{3}$, despite forming a generating set of $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$.
Case $(d,e)=(2,2)$: by arguments analogous to those used in the previous case, we obtain $X^{2},Y^{2}\in\mathfrak{a}\cap\mathfrak{b}$. Since $\langle X^{2},Y^{2}\rangle^{2}=\langle X^{4},X^{2}Y^{2},Y^{4}\rangle\subsetneq\mathfrak{c}_{4}$, we may assume without loss of generality that $\langle X^{2},Y^{2}\rangle\subsetneq\mathfrak{a}$. By Lemma 4.1, there exist nonnegative integers $i$ and $j$ such that $X^{i}Y^{j}\in\mathfrak{a}$, and neither $X^{2}$ nor $Y^{2}$ divides $X^{i}Y^{j}$; since $\operatorname{\mathsf{m-deg}}(X^{i}Y^{j})\geq d=2$, the only possibility is $(i,j)=(1,1)$, i.e., $XY\in\mathfrak{a}$. Multiplying with $Y^{2}\in\mathfrak{b}$, we obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, a contradiction (as in the previous case).
@@
Remark 4.4.
An important consequence of Lemma 4.1 is that every monomial ideal II admits a unique generating set of monomials that is minimal with respect to divisibility [13, Prop. 1.11]; we denote this set by $G(I)$. As a consequence of Dickson’s Lemma [13, Theorem 1.9], $G(I)$ is finite for every $I\in{\mathcal{M}{\rm on}}(R)$.
Next, we see to what extent can [21, Theorem 5.1] be adapted to the monoid ${\mathcal{M}{\rm on}}(R)$. Let us define a further class of monomial ideals, namely $\mathfrak{a}_{k}={\langle X,Y\rangle}^{k}$, for any $k\in\mathbb{N}^{+}$.
@@
(1) The fact that ${\mathcal{M}{\rm on}}(R)$ is not locally finitely generated follows from the proof of the corresponding statement for ${\mathcal{I}}(R)$ in [21, Lemma 5.3.2], since all the ideals involved are monomial, and a monomial ideal which is an atom in ${\mathcal{I}}(R)$ is clearly also an atom in ${\mathcal{M}{\rm on}}(R)$. For the second part of the statement, assume to the contrary that there are a Krull monoid $H$ and a transfer homomorphism $\theta\colon{\mathcal{M}{\rm on}}(R)\to H$. Consider the monomial ideals
@@
A straightforward computation shows that $\mathfrak{a}_{1}\mathfrak{a}_{2}=\mathfrak{a}_{1}\mathfrak{b}_{2}$, whence $\theta(\mathfrak{a}_{1})\theta(\mathfrak{a}_{2})=\theta(\mathfrak{a}_{1})\theta(\mathfrak{b}_{2})$ in $H$. Since Krull monoids are cancellative, we infer $\theta(\mathfrak{a}_{2})=\theta(\mathfrak{b}_{2})$. By Lemma 2.1, for any $I\in{\mathcal{M}{\rm on}}(R)$ we have that $I\in\mathcal{A}({\mathcal{M}{\rm on}}(R))$ if and only if $\theta(I)\in\mathcal{A}(H)$. Since $\mathfrak{b}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, it follows that $\theta(\mathfrak{b}_{2})$ is an atom of $H$. Thus, $\theta(\mathfrak{a}_{2})$ is also an atom of $H$, whence $\mathfrak{a}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, a contradiction.
(2) One can proceed exactly as in [21, Lemma 5.3.3] to show that $\mathcal{U}_{2}({\mathcal{M}{\rm on}}(R))=\mathbb{N}_{\geq 2}$, as all the ideals occurring in that proof are monomial. Since ${\mathcal{M}{\rm on}}(R)$ is a BF-monoid, the conclusion then follows by [21, Lemma 3.1.3].
@@
involving only monomial ideals, and recalling Remark 4.3, we can conclude that $2\in L_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{5})$.
(4) Set $\mathfrak{p}=\langle X\rangle$, denote by $H_{1}$ the free monoid with basis $\{\mathfrak{p}\}$ and let $H_{2}=\{\mathfrak{a}\in{\mathcal{M}{\rm on}}(R)\colon\mathfrak{p}\nmid\mathfrak{a}\}$. The ideal $\mathfrak{p}$ is principal, thus invertible in the monoid of nonzero fractional ideals of $R$. In particular, it is a cancellative element of this monoid, hence also a cancellative element of ${\mathcal{M}{\rm on}}(R)$. Moreover, $\mathfrak{p}$ is clearly prime, whence we obtain ${\mathcal{M}{\rm on}}(R)=H_{1}\times H_{2}$. Now observe that $\mathfrak{a}_{k}\in H_{2}$ for all $k\in\mathbb{N}^{+}$, as $Y^{k}\in\mathfrak{a}_{k}$. By item (3), we can then apply [21, Prop. 3.2] to conclude that ${\mathcal{M}{\rm on}}(R)$ is fully elastic. ∎
@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
We continue our investigation of ${\mathcal{M}{\rm on}}(R)$ by constructing new classes of atoms in this monoid. Fix a positive integer $n\geq 3$ and $a_{1},\dots,a_{n},a_{n+1}\in\mathbb{N}^{+}$ verifying conditions (C1), $a_{n+1}=a_{1}+\dots+a_{n-1}+2a_{n}$, and (C2), $a_{i+1}>2(a_{1}+\dots+a_{i})$ for all $i\in\llbracket 1,n-1\rrbracket$. Keeping the notation of Section 3.2, for every nonempty $I\subseteq\llbracket 1,n+1\rrbracket$ we set $a_{I}=\sum_{i\in I}a_{i}$ while we set $a_{\emptyset}=0$. For $r\in\llbracket 3,n\rrbracket$ let us define:
@@
Thus, if there exists $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ such that $X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}$ divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, then it must be $a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}\leq a_{\llbracket 3,r\rrbracket}-a_{2}$, which implies $a_{J}\geq a_{1}+a_{2}$ and, by (9), $a_{J}\geq a_{3}$. Moreover, we must also have $a_{J}\leq a_{3}-a_{2}$, but this contradicts the previous inequality.
@@
The construction of the class of atoms $\tilde{\mathfrak{b}}_{r}$ of ${\mathcal{M}{\rm on}}(R)$ gives us the opportunity to answer negatively, for this monoid, the question posed ad the end of Section 3.
@@
We already know by Theorem 3.7(2) that $\{2,n+1\}\subseteq\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})\subseteq\llbracket 2,n+1\rrbracket$ for every $n\geq 2$. If $n=2$ the conclusion is trivial; if $n\geq 3$ it follows from Proposition 4.6(6) and Theorem 4.7. ∎
@@
We first observe that $a_{J}-a_{I}=a_{J\setminus I}-a_{I\setminus J}$. If $I\setminus J=\emptyset$, then $a_{J}-a_{I}=a_{J\setminus I}$ and $a_{J\setminus I}>a_{1}$ trivially implies $a_{J\setminus I}\geq a_{3}$, as $J\setminus I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$. Assume then $I\setminus J\neq\emptyset$. Clearly our assumptions imply that also $J\setminus I\neq\emptyset$ (otherwise $a_{J}-a_{I}<0$), so there exist $s,t>0$ such that $s=\max(J\setminus I)$ and $t=\max(I\setminus J)$. Note that, since $(J\setminus I)\cap(I\setminus J)=\emptyset$, $s\neq t$. Moreover, it must be $s>t$. In fact, if $s<t$ (implying that $t\geq 3$), then
@@
In the following proofs, we will use without mention the following property of minimal generating sets of products of monomial ideals, whose proof is straightforward: given $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$, we have
$G(\mathfrak{a}\cdot\mathfrak{b})\subseteq\{fg\colon f\in G(\mathfrak{a}),g\in G(\mathfrak{b})\}.$
@@
(8) First we prove that $X^{d},Y^{d}\in\mathfrak{a}$ and $X^{e},Y^{e}\in\mathfrak{b}$. By equation (1), $m=d+e$. Now if $X^{i}\not\in\mathfrak{a}$ for every $i\in\llbracket 1,m\rrbracket$, then $X^{m}\not\in I$. Therefore $X^{i_{0}}\in\mathfrak{a}$ for some $i_{0}\in\llbracket 1,m\rrbracket$ such that $i_{0}\geq d$ and hence $X^{m-i_{0}}\in\mathfrak{b}$, where $m-i_{0}\geq e$. If $i_{0}>d$, then $i_{0}+e>d+e=m$ and hence $e>m-i_{0}$, which contradicts the fact that $\operatorname{\mathsf{m-deg}}{(\mathfrak{b})}=e$. So $i_{0}=d$ and hence $X^{d}\in\mathfrak{a}$ and $X^{e}\in\mathfrak{b}$. Likewise, we can show that $Y^{d}\in\mathfrak{a}$ and $Y^{e}\in\mathfrak{b}$. Since $d=\operatorname{\mathsf{m-deg}}{(\mathfrak{a})}$, $X^{d},Y^{d}\in\mathfrak{a}$ and $\mathfrak{a}$ is a monomial ideal, by Lemma 4.1 we can conclude that $X^{d},Y^{d}\in G(\mathfrak{a})$; analogously, $X^{e},Y^{e}\in G(\mathfrak{b})$. ∎
@@
(9) Follows from Equation (1) and Lemma 4.11(7).
@@
By (10) $X^{\alpha},Y^{\alpha}\in\mathfrak{a}_{K}[\alpha]$ and $X^{\beta},Y^{\beta}\in\mathfrak{b}_{K}[\beta]$. Hence, $X^{\alpha}Y^{\beta}\in\tilde{\mathfrak{b}}_{r,K}[a_{\{1\}\cup\llbracket 3,r\rrbracket}]$, and so

### v2

@@ 4. Arithmetical properties of ${\mathcal{M}{\rm on}}(R)$
Let $I\in{\mathcal{M}{\rm on}}(R)$, and $\mathcal{G}$ a set of monomials in $I$. Then $\mathcal{G}$ generates $I$ if and only if for each monomial $v\in I$ there exists $u\in\mathcal{G}$ such that $u\mid v$.
Remark 4.2.
If $\mathcal{G}$ is a set of monomials, denote by $\operatorname{\mathrm{Min}}\mathcal{G}$ the subset consisting of those monomials that are minimal with respect to divisibility among the elements of $\mathcal{G}$. An important consequence of Lemma 4.1 is that every monomial ideal $I$ admits a unique generating set of monomials that is minimal with respect to inclusion [13, Prop. 1.11]. More precisely, if $\mathcal{G}$ is any set of monomials such that $I=\langle\mathcal{G}\rangle$, then $\operatorname{\mathrm{Min}}\mathcal{G}$ is the minimal generating set of $I$; we denote it by $G(I)$. As a consequence of Dickson’s Lemma [13, Theorem 1.9], $G(I)$ is finite for every $I\in{\mathcal{M}{\rm on}}(R)$.
The following corollary of Lemma 4.1 will be used throughout the section: in essence, it tells us that when an ideal generated only by monomials in $X$ and $Y$ factors into a product of monomial ideals, the factors can also be taken to have only monomials in $X$ and $Y$ as generators. In preparation for its proof, we isolate the following special case of Remark 4.2: if $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}{(R)}$, then $G(\mathfrak{a}\cdot\mathfrak{b})={\operatorname{\mathrm{Min}}\{fg\colon f\in G(\mathfrak{a}),\ g\in G(\mathfrak{b})\}}$.
Corollary 4.3.
Suppose $I=\mathfrak{a}\cdot\mathfrak{b}$ for some $I,\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)$ and $G(I)\subseteq K[X,Y]$. Let $\mathfrak{a}^{\prime}$ and $\mathfrak{b}^{\prime}$ be the ideals generated, respectively, by the sets of monomials $G(\mathfrak{a})\cap K[X,Y]$ and $G(\mathfrak{b})\cap K[X,Y]$. Then $I=\mathfrak{a}^{\prime}\cdot\mathfrak{b}^{\prime}$.
Proof.
Suppose that there exists $u\in G(\mathfrak{a})\setminus K[X,Y]$. It follows that $X_{k}$ divides $u$ for some $k\in\llbracket 3,N\rrbracket$, whence $X_{k}$ also divides every monomial in $\{uv\colon v\in G(\mathfrak{b})\}$. Since the monomials in $G(\mathfrak{a}\cdot\mathfrak{b})=G(I)$ are not divisible by $X_{k}$, we have $uv\not\in\operatorname{\mathrm{Min}}\{fg\colon f\in G(\mathfrak{a}),\ g\in G(\mathfrak{b})\}$ for every $v\in G(\mathfrak{b})$, whence
$I$ $=\mathfrak{a}\cdot\mathfrak{b}$
$=\langle G(\mathfrak{a}\cdot\mathfrak{b})\rangle$
$=\langle\operatorname{\mathrm{Min}}\{fg\colon f\in G(\mathfrak{a}),\ g\in G(\mathfrak{b})\}\rangle$
$=\langle\operatorname{\mathrm{Min}}\{fg\colon f\in G(\mathfrak{a})\setminus\{u\},\ g\in G(\mathfrak{b})\}\rangle$
$=\langle G(\mathfrak{a})\setminus\{u\}\rangle\cdot\langle G(\mathfrak{b})\rangle\text{.}$
Repeating this argument for every $u\in G(\mathfrak{a})\setminus K[X,Y]$ and every $v\in G(\mathfrak{b})\setminus K[X,Y]$ yields the desired conclusion (note that $G(\mathfrak{a})$ and $G(\mathfrak{b})$ are finite sets, by Remark 4.2). ∎
@@
is an atom of ${\mathcal{M}{\rm on}}(R)$, while back in Section 2 we recalled that it is not an atom of ${\mathcal{I}}(R)$ (when the characteristic of $K$ is not $2$), despite $\{0,1,2,4\}$ being an atom of $\mathcal{P}_{\rm fin}(\mathbb{N})$. Suppose that $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$ for some $\mathfrak{a},\mathfrak{b}\in{\mathcal{M}{\rm on}}(R)\setminus\{R\}$, with $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$ and $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$. By Corollary 4.3, we can assume that all the generators of $\mathfrak{a}$ and $\mathfrak{b}$ are monomials in $X$ and $Y$. By equation (1) and Remark 4.4, $d+e=4$, $d\geq 1$ and $e\geq 1$. By symmetry, we can also assume $d\leq e$, so the only possibilities for $(d,e)$ are $(1,3)$ and $(2,2)$.
Case $(d,e)=(1,3)$: first, we claim that $Y^{3}\in G(\mathfrak{b})$. Indeed, if this is not the case, for every monomial $X^{i}Y^{j}\in G(\mathfrak{b})$ we have either $i\geq 1$ or $j\geq 4$, as $e=3$; the product of such elements with monomials of $G(\mathfrak{a})$ yields elements of the form $X^{k}Y^{\ell}$, with either $k\geq 1$ or $\ell\geq 5$. In particular, $Y^{4}\not\in G(\mathfrak{a}\cdot\mathfrak{b})$, a contradiction. Now suppose $X\notin G(\mathfrak{a})$; since $d=1$, for every monomial $X^{i}Y^{j}\in G(\mathfrak{a})$ we have either $i\geq 2$ or $j\geq 1$. The product of such elements with monomials of $G(\mathfrak{b})$ yields elements of the form $X^{k}Y^{\ell}$, with either $k+\ell\geq 5$ or $\ell\geq 1$. In particular, $X^{4}\not\in G(\mathfrak{a}\cdot\mathfrak{b})$, again a contradiction; therefore, $X\in G(\mathfrak{a})$. We then obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, absurd by Lemma 4.1, as none of the monomials $X^{4},X^{3}Y,X^{2}Y^{2},Y^{4}$ divide $XY^{3}$, despite forming a generating set of $\mathfrak{c}_{4}=\mathfrak{a}\cdot\mathfrak{b}$.
Case $(d,e)=(2,2)$: arguing in the previous case, we obtain $X^{2},Y^{2}\in G(\mathfrak{a})\cap G(\mathfrak{b})$. Since $\langle X^{2},Y^{2}\rangle^{2}=\langle X^{4},X^{2}Y^{2},Y^{4}\rangle\subsetneq\mathfrak{c}_{4}$, we may assume without loss of generality that $\langle X^{2},Y^{2}\rangle\subsetneq\mathfrak{a}$. By Lemma 4.1, there exist nonnegative integers $i$ and $j$ such that $X^{i}Y^{j}\in G(\mathfrak{a})$, and neither $X^{2}$ nor $Y^{2}$ divides $X^{i}Y^{j}$; since $\operatorname{\mathsf{m-deg}}(X^{i}Y^{j})\geq d=2$, the only possibility is $(i,j)=(1,1)$, i.e., $XY\in G(\mathfrak{a})$. Multiplying with $Y^{2}\in\mathfrak{b}$, we obtain $XY^{3}\in\mathfrak{a}\cdot\mathfrak{b}$, a contradiction (as in the previous case).
@@
Next, we see to what extent [21, Theorem 5.1] can be adapted to the monoid ${\mathcal{M}{\rm on}}(R)$. Let us define a further class of monomial ideals, namely $\mathfrak{a}_{k}={\langle X,Y\rangle}^{k}$, for any $k\in\mathbb{N}^{+}$.
@@
(1) The fact that ${\mathcal{M}{\rm on}}(R)$ is not locally finitely generated follows from the proof of the corresponding statement for ${\mathcal{I}}(R)$ in [21, Lemma 5.3.2], since all the ideals involved are monomial, and a monomial ideal that is an atom in ${\mathcal{I}}(R)$ is clearly also an atom in ${\mathcal{M}{\rm on}}(R)$. For the second part of the statement, assume to the contrary that there are a Krull monoid $H$ and a transfer homomorphism $\theta\colon{\mathcal{M}{\rm on}}(R)\to H$. Consider the monomial ideals
@@
A straightforward computation shows that $\mathfrak{a}_{1}\mathfrak{a}_{2}=\mathfrak{a}_{1}\mathfrak{b}_{2}$, whence $\theta(\mathfrak{a}_{1})\theta(\mathfrak{a}_{2})=\theta(\mathfrak{a}_{1})\theta(\mathfrak{b}_{2})$ in $H$. Since Krull monoids are cancellative, we infer $\theta(\mathfrak{a}_{2})=\theta(\mathfrak{b}_{2})$. By Lemma 2.1, for any $I\in{\mathcal{M}{\rm on}}(R)$ we have $I\in\mathscr{A}({\mathcal{M}{\rm on}}(R))$ if and only if $\theta(I)\in\mathscr{A}(H)$. Since $\mathfrak{b}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, it follows that $\theta(\mathfrak{b}_{2})$ is an atom of $H$. Thus, $\theta(\mathfrak{a}_{2})$ is also an atom of $H$, whence $\mathfrak{a}_{2}$ is an atom of ${\mathcal{M}{\rm on}}(R)$, a contradiction.
(2) One can proceed exactly as in [21, Lemma 5.3.3] to show that $\mathcal{U}_{2}({\mathcal{M}{\rm on}}(R))=\mathbb{N}_{\geq 2}$, as all the ideals occurring in that proof are monomial. Since ${\mathcal{M}{\rm on}}(R)$ is a BF-monoid, the conclusion follows by [21, Lemma 3.1.3].
@@
involving only monomial ideals, and recalling Example 4.5, we can conclude that $2\in L_{{\mathcal{M}{\rm on}}(R)}(\mathfrak{a}_{5})$.
(4) Set $\mathfrak{p}=\langle X\rangle$, denote by $H_{1}$ the free monoid with basis $\{\mathfrak{p}\}$ and let $H_{2}=\{\mathfrak{a}\in{\mathcal{M}{\rm on}}(R)\colon\mathfrak{p}\nmid\mathfrak{a}\}$. The ideal $\mathfrak{p}$ is principal and therefore invertible in the monoid of nonzero fractional ideals of $R$. In particular, it is a cancellative element of this monoid, and hence also a cancellative element of ${\mathcal{M}{\rm on}}(R)$. Moreover, $\mathfrak{p}$ is clearly prime, whence we obtain ${\mathcal{M}{\rm on}}(R)=H_{1}\times H_{2}$. Now observe that $\mathfrak{a}_{k}\in H_{2}$ for all $k\in\mathbb{N}^{+}$, as $Y^{k}\in\mathfrak{a}_{k}$. By item (3), we can then apply [21, Prop. 3.2] to conclude that ${\mathcal{M}{\rm on}}(R)$ is fully elastic. ∎
@@ 4.1. Atoms and sets of lengths in ${\mathcal{M}{\rm on}}(R)$
We continue our investigation of ${\mathcal{M}{\rm on}}(R)$ by constructing new classes of atoms in this monoid. Fix a positive integer $n\geq 3$ and $a_{1},\dots,a_{n},a_{n+1}\in\mathbb{N}^{+}$ verifying conditions (C1), $a_{n+1}=a_{1}+\dots+a_{n-1}+2a_{n}$, and (C2), $a_{i+1}>2(a_{1}+\dots+a_{i})$ for all $i\in\llbracket 1,n-1\rrbracket$. Keeping the notation of Section 3.2, for every nonempty $I\subseteq\llbracket 1,n+1\rrbracket$ we set $a_{I}=\sum_{i\in I}a_{i}$, while setting $a_{\emptyset}=0$. For $r\in\llbracket 3,n\rrbracket$ let us define:
@@
Thus, if there exists $J\subseteq\{1\}\cup\llbracket 3,r\rrbracket$ such that $X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}$ divides $X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}$, then it must be $a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}\leq a_{\llbracket 3,r\rrbracket}-a_{2}$, which implies $a_{J}\geq a_{1}+a_{2}$ and, by (9), $a_{J}\geq a_{3}$. Moreover, we must also have $a_{J}\leq a_{3}-a_{2}$, but this contradicts the previous inequality. We can conclude that
$G(\tilde{\mathfrak{b}}_{r})=\{X^{a_{1}+a_{\llbracket 3,r\rrbracket}-a_{J}}Y^{a_{J}}\colon J\subseteq\{1\}\cup\llbracket 3,r\rrbracket\}\cup\{X^{a_{\llbracket 3,r\rrbracket}-a_{2}}Y^{a_{3}-a_{2}}\}.$
@@
The construction of the class of atoms $\tilde{\mathfrak{b}}_{r}$ of ${\mathcal{M}{\rm on}}(R)$ gives us the opportunity to answer negatively, for this monoid, the question posed at the end of Section 3.
@@
Fix $n\geq 2$. By Remark 3.6, we obtain $\{2,n+1\}\subseteq\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})$, as the factorizations therein are into monomial ideals which are known to be atoms of ${\mathcal{I}}(R)$, and so also of ${\mathcal{M}{\rm on}}(R)$. By Theorem 3.7(2), we have $\mathsf{L}_{{\mathcal{I}}(R)}(I_{C_{n}})\subseteq\llbracket 2,n+1\rrbracket$. Since any factorization $z\in\mathsf{Z}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})$ can be expanded into a factorization $w\in\mathsf{Z}_{{\mathcal{I}}(R)}(I_{C_{n}})$ with $\lvert w\rvert\geq\lvert z\rvert$, we obtain $\sup\mathsf{L}_{{\mathcal{I}}(R)}(I_{C_{n}})\geq\sup\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})$, whence $\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})$ is also contained in $\llbracket 2,n+1\rrbracket$ ($1\not\in\mathsf{L}_{{\mathcal{M}{\rm on}}(R)}(I_{C_{n}})$ as $I_{C_{n}}$ is not an atom). If $n=2$, the conclusion is trivial; if $n\geq 3$, it follows from Proposition 4.7(6) and Theorem 4.8. ∎
@@
We first observe that $a_{J}-a_{I}=a_{J\setminus I}-a_{I\setminus J}$. If $I\setminus J=\emptyset$, then $a_{J}-a_{I}=a_{J\setminus I}$ and $a_{J\setminus I}>a_{1}$ trivially implies $a_{J\setminus I}\geq a_{3}$, as $J\setminus I\subseteq\{1\}\cup\llbracket 3,r\rrbracket$. Assume then $I\setminus J\neq\emptyset$. Clearly, our assumptions also imply $J\setminus I\neq\emptyset$ (otherwise $a_{J}-a_{I}<0$), so there exist $s,t>0$ such that $s=\max(J\setminus I)$ and $t=\max(I\setminus J)$. Note that, since $(J\setminus I)\cap(I\setminus J)=\emptyset$, $s\neq t$. Moreover, it must be $s>t$. In fact, if $s<t$ (implying that $t\geq 3$), then
@@
(8) Since $m=\operatorname{\mathsf{m-deg}}(I)$ and $X^{m},Y^{m}\in I$, by Lemma 4.1 we deduce $X^{m},Y^{m}\in G(I)$. As $G(I)=G(\mathfrak{a}\cdot\mathfrak{b})\subseteq\{fg\colon f\in G(\mathfrak{a}),g\in G(\mathfrak{b})\}$ and $X^{m}\in G(I)$, necessarily $X^{i}\in G(\mathfrak{a})$ and $X^{j}\in G(\mathfrak{b})$ for some $i,j$ such that $i+j=m$. Since $\operatorname{\mathsf{m-deg}}(\mathfrak{a})=d$ and $\operatorname{\mathsf{m-deg}}(\mathfrak{b})=e$, we have $i\geq d$ and $j\geq e$, whence equation (1) (i.e., $m=d+e$) implies $i=d$ and $j=e$. Analogously, we can prove $Y^{d}\in G(\mathfrak{a})$ and $Y^{e}\in G(\mathfrak{b})$. ∎
@@
(9) Follows from Lemma 4.12(7) and equation (1).
@@
By (10), $X^{\alpha},Y^{\alpha}\in\mathfrak{a}_{K}[\alpha]$ and $X^{\beta},Y^{\beta}\in\mathfrak{b}_{K}[\beta]$. Hence, $X^{\alpha}Y^{\beta}\in{(\tilde{\mathfrak{b}}_{r})}_{K}[a_{\{1\}\cup\llbracket 3,r\rrbracket}]$, and so